OpenStudy (anonymous):

Suppose the point (2, 4) is on the graph of y = f(x). Find a point on the graph of the given function. y = f(x + 1)

OpenStudy (jamesj):

Well, we know that 4 = f(2) and that all we know about f. So what are the only values of x and y can we use in y = f(x+1) ?

OpenStudy (anonymous):

A) (1, 4) B) (2, 5) C) (3, 4) or D) (2, 3)

OpenStudy (jamesj):

Forget the possible answers now and think. We only know ONE thing about this function f(x): that (x,y) = (2,4) is on the graph. So now that we need to find a point on the graph of y = f(x+1) what can be the ONLY choice for y?

OpenStudy (jamesj):

y = f(x+1) and 4 = f(2), hence y must be .... what?

OpenStudy (anonymous):

4

OpenStudy (jamesj):

Yep. Now what must x be?

OpenStudy (jamesj):

(mathematical eagles: the "must" here isn't quite right, but for this question it's fine.)

myininaya (myininaya):

y=f(x+1) has ordered pairs of the form (x+1,y) y=f(x) has ordered pairs of the form (x,y) we know a point on y=f(x) which is (2,4) so when does (2,4)=(x+1,y) you said y=4 so we have (2,4)=(x+1,4)